OpenStudy (anonymous):

Hockey team receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 64 points. They won 11 more games than they tied. How many wins and how many ties did the team have? how many ties?

OpenStudy (anonymous):

okay...where are you at?

OpenStudy (anonymous):

I know that there are 25 wins..

OpenStudy (anonymous):

okay...how do you know that?

OpenStudy (anonymous):

2(y+11)+y=64 3y+22=64 y=14 x=14+11=25

OpenStudy (anonymous):

okay....good...so you got your equations down...so you just want to know the number of games tied...remember you set y = number of games tied correct? you also said that x = y+11 since you now have x...can you plug that in and solve for y?

OpenStudy (anonymous):

so 36

OpenStudy (anonymous):

how bout 14 instead

OpenStudy (anonymous):

14 is correct...please check your answer